### v1

# Pretraining Scaling Laws for Generative Evaluations of Language Models

Abstract
Neural scaling laws have played a central role in modern machine learning, driving the field’s ever-expanding scaling of parameters, data and compute. While much research has gone into fitting scaling laws and predicting performance on pretraining losses and on discriminative evaluations such as multiple-choice question-answering benchmarks, comparatively little research has been done on fitting scaling laws and predicting performance on generative evaluations such as mathematical problem-solving or coding. In this work, we propose and evaluate three different pretraining scaling laws for fitting pass-at-$k$ on generative evaluations and for predicting pass-at-$k$ of the most expensive model using the performance of cheaper models. Our three scaling laws differ in the covariates used: (1) pretraining compute, (2) model parameters and pretraining tokens, (3) log likelihoods of gold reference solutions. We make four main contributions: First, we show how generative evaluations offer new hyperparameters (in our setting, $k$) that researchers can use to control the scaling laws parameters and the predictability of performance. Second, in terms of scaling law parameters, we find that the compute scaling law and parameters + tokens scaling law stabilize for the last $\mathord{\sim}1.5{-}2.5$ orders of magnitude, whereas the gold reference likelihood scaling law stabilizes for the last $\mathord{\sim}5$ orders of magnitude. Third, in terms of predictive performance, we find all three scaling laws perform comparably, although the compute scaling law predicts slightly worse for small $k$ and the log likelihoods of gold reference solutions predicts slightly worse for large $k$. Fourth, we establish a theoretical connection that the compute scaling law emerges as the compute-optimal envelope of the parameters-and-tokens scaling law. Our framework provides researchers and practitioners with insights and methodologies to forecast generative performance, accelerating progress toward models that can reason, solve, and create.
1 Introduction
Neural scaling laws, which predictably map resources to the performance of neural networks, are foundational for developing frontier AI systems (Kaplan et al., 2020b; Hoffmann et al., 2022b). Such empirical regularities for improving model performance from scaling parameters, data, and compute have driven pursuit of ever-larger models. Much of the research in neural scaling laws has focused on fitting and predicting model performance on pretraining losses, e.g., (Kaplan et al., 2020b; Hoffmann et al., 2022b; Clark et al., 2022; Hernandez et al., 2022; Sardana et al., 2024; Muennighoff et al., 2023; Porian et al., 2024; Pearce & Song, 2024) and on “downstream” discriminative evaluations like multiple-choice question-answering, e.g., (Schaeffer et al., 2023; Gadre et al., 2024; Schaeffer et al., 2024; Grattafiori et al., 2024; Chen et al., 2025; Bhagia et al., 2025).
While both pretraining losses and discriminative evaluations are undeniably useful, many capabilities we care about are generative: writing books, autoformalizing mathematics, or conducting cutting-edge research. Comparatively little research has investigated scaling laws for generative evaluations (OpenAI et al., 2024; Hu et al., 2024); for a deeper discussion of related work, please see Appendix A. Unlike pretraining losses and discriminative evaluations, generative evaluations introduce new considerations such as which metric is used to quantify performance and how samples are drawn from the model, including the sampling temperature (Ackley et al., 1985) and the sampling algorithm, e.g., top-k (Fan et al., 2018), top-p (Holtzman et al., 2020).
In this work, we propose a framework for fitting scaling laws and predicting performance on generative evaluations, focusing on a particular setting: benchmarks with verifiable binary rewards, with multiple attempts per problem, with performance scored using the “pass-at-kk” metric (Kulal et al., 2019; Chen et al., 2021). We chose this setting due to its widespread use in the literature (First et al., 2023; Brown et al., 2024; Hassid et al., 2024; Hughes et al., 2024; Chen et al., 2024; Ehrlich et al., 2025; Kwok et al., 2025). In this setting, we make the following contributions:
  1. 1.
    We propose three scaling laws that fit and predict pass-at-$k$ from three different covariates: (1) pretraining compute (Sec. 3), (2) model parameters and pretraining tokens (Sec. 4) and (3) likelihoods of gold reference solutions (Sec. 5).
  2. 2.
    We reveal that a hyperparameter specific to this setting – the number of attempts per problem $k$ – offers a new lever to change the parameters of the scaling laws as well as how accurately the scaling laws extrapolate to more expensive models.
  3. 3.
    We quantify how stable each scaling law’s parameters are, finding that parameters for the compute scaling law and the parameters + tokens scaling law stabilize by the last $\mathord{\sim}1.5{-}2.5$ orders of magnitude of pretraining compute, whereas parameters for the gold reference likelihood scaling law are stable over the last $\mathord{\sim}5$ orders of magnitude.
  4. 4.
    We quantify how predictive each scaling law is, finding that all three comparably predict the performance of the most expensive model based on cheaper models. As a lesser difference, we find that the compute scaling law has slightly higher predictive error for small $k$, whereas the gold reference likelihoods law has slightly higher predictive error for large $k$.
  5. 5.
    We prove that the compute-only scaling law is the compute-optimal envelope of the parameters-and-tokens scaling law, obtained by minimizing the objective under a fixed compute budget. We also derive the optimal split of compute between model parameters and pretraining data, and a dimensionless misallocation penalty that explains when fixed training recipes underperform compute-optimal scaling.
Figure 1: Scaling of Pass Rates with Pretraining Compute (Full Fit). Each panel fits Eqn. 3 $\log\!\big(\mathrm{pass}_{\mathcal{B}}@k\big)(C,k)=E_{0}(k)+C_{0}(k)\cdot C^{-\alpha(k)}$ to GSM8K pass rates for Pythia checkpoints across $\sim$5 orders of magnitude of pretraining compute. Panels show $k\in\{1,10^{2},10^{4}\}$. The fit curves capture the trend moderately well and reveal two effects of increasing $k$: (i) the irreducible error $\widehat{E}_{0}(k)$ falls from $2\to 0$ as $k$ increases from $1\to 100$; and (ii) the compute-dependent terms increase with $k$: the prefactor grows from $\widehat{C}_{0}(k=1)\approx 1e3$ to $\widehat{C}_{0}(k=10000)\!\sim\!7.8e5$ and the exponent rises from $\widehat{\alpha}(k=1)\!\approx\!0.120$ to $\widehat{\alpha}(k=100)\!\approx\!0.325$. Takeaway: Larger $k$ makes the relationship a pure power law and makes pass rates a steeper function of pretraining compute.
2 Methodology
Benchmark To study how language model performance changes during pretraining on generative evaluations, we used GSM8K (Cobbe et al., 2021), a well-established benchmark containing $8,500$ grade school math word problems designed to evaluate a model’s ability to perform multi-step mathematical reasoning. Each attempt at solving a problem yields a binary outcome: the attempt is either successful or unsuccessful. Due to limited budget, we used a randomly chosen subset of $128$ problems. See Appendix B for a description of how many samples we drew per problem per model.
Metric To evaluate the performance of each model, we used the ubiquitous pass-at-$k$ metric (Kulal et al., 2019). For the $i$-th problem, a model’s pass rate at $k$ attempts is the probability of the event that any of $k$ attempts are successful. Because generative evaluations sample from the language model, this per-problem pass rate must be estimated. We use the unbiased, low variance and numerically stable estimator introduced by Chen et al. (2021), which, for each problem, draws $n_{i}>k$ samples, counts the number of successes $s_{i}\in[0,n_{i}]$, and then averages over all subsets of size $k$:
$\mathrm{pass_{i}@k}\quad\operatorname{\stackrel{{\scriptstyle\text{def}}}{{\;=\;}}}\quad 1-\dbinom{n_{i}-s_{i}}{k}/\dbinom{n_{i}}{k}$. (1)
For a benchmark $\mathcal{B}$ of $|\mathcal{B}|$ problems, a model’s benchmark pass rate at $k$ attempts is estimated as:
$\mathrm{pass_{\mathcal{B}}@k}\quad\operatorname{\stackrel{{\scriptstyle\text{def}}}{{\;=\;}}}\quad\frac{1}{|\mathcal{B}|}\,\sum_{i\in\mathcal{B}}\,\mathrm{pass_{i}@k}$ (2)
Pass rates are sometimes phrased as the fraction of problems in the benchmark solved (“coverage”), but these two quantities are mathematically equivalent (Brown et al., 2024; Schaeffer et al., 2025b).
Figure 2: Scaling Law Parameters Vary with Attempts per Problem (Full Fit). In comparison with pretraining loss scaling laws and discriminative evaluation scaling laws, the number of attempts per problem $k$ introduces a new hyperparameter that affects the parameters of generative evaluation scaling laws, and also how predictable $\mathrm{pass_{\mathcal{B}}}@k$ is. Here we show the effect of $k$ for the pretraining compute scaling law (Eqn. 3); for a side-by-side comparison with our next two scaling laws, see Appendix C. Left: The irreducible error $E_{0}(k)$ falls roughly exponentially with $k$ and is $\approx 0$ by $k\!\approx\!10^{2}$, consistent with the probability that a problem remains unsolved decaying exponentially in the number of attempts. Center: The compute prefactor $C_{0}(k)$ increases monotonically with $k$, indicating that once $E_{0}(k)\!\to\!0$, variation is dominated by the compute-dependent term. Right: The compute exponent $\alpha(k)$ increases smoothly, from $\sim\!0.15$ at k=1k{=}1 to $\sim\!0.3$ at $k{=}10^{4}$, implying steeper and more regular scaling at larger $k$. Hue corresponds to $k$.
Sampling We used temperature-only sampling (Ackley et al., 1985) at $\tau{=}1.0$ and did not sweep on account of prior work (Schaeffer et al., 2025a).
Language Model Family For our experiments, we used the Pythia family (Biderman et al., 2023), which contains 8 models from 14M to 12B parameters pretrained on 300B tokens from The Pile (Gao et al., 2020). We used 8 checkpoints per parameter size of the non-deduplicated variants, but excluded extremely overtrained checkpoints identified as abnormal by prior work (Godey et al., 2024). Each model checkpoint has a specific number of model parameters ($N$) and has been trained on a certain number of tokens ($D$). We took the standard approximation for pretraining compute of $C\approx 6\cdot N\cdot D$ (Kaplan et al., 2020b; Hoffmann et al., 2022b; Pearce & Song, 2024; Gadre et al., 2024; Schaeffer et al., 2024; Porian et al., 2024). We use intermediate checkpoints in lieu of fully trained models, and note that our results would likely improve with greater experimental control.
3 Fitting and Predicting Pass Rates from Pretraining Compute
For our first scaling law, we assessed how well the benchmark $\mathrm{pass_{\mathcal{B}}@k}$ can be (1) fit as a function of pretraining compute $C$ and (2) predicted from pretraining compute.
3.1 Fitting Pass Rates from Pretraining Compute
Motivated by the GPT-4 Technical Report (OpenAI et al., 2024), we posited that the negative log of benchmark pass rates follows a scaling law as a function of model pretraining compute $C$:
$\log\big(\mathrm{pass_{\mathcal{B}}}@k\big)(C,k)=E_{0}(k)+\frac{C_{0}(k)}{C^{\alpha(k)}},$ (3)
where $E_{0}(k)\geq 0$ is the irreducible error, $C_{0}(k)>0$ is the compute scaling prefactor, and $\alpha(k)>0$ is the compute scaling exponent. In contrast with prior research on scaling laws, parameters that were previously constant are now parameterized by the number of attempts per problem $k$. Thus, $k$ can be viewed as a hyperparameter that offers a new lever to change the predictability of scaling laws in a way that is not available for pretraining loss or discriminative evaluation scaling laws.
Figure 3: Predicting Pass Rates from Scaling Pretraining Compute (Backtesting). We evaluate predictability via backtesting: we iteratively fit Eq. 3 on subsets of models up to a maximum pretraining compute, then extrapolate to predict the most expensive model’s $\log\!\big(\mathrm{pass}_{\mathcal{B}}@k\big)$. The most expensive model is Pythia 12B-parameter 300B-token ($\approx 2.16\times 10^{22}$ FLOP). The $x$-axis in all panels is the ratio of the largest-compute checkpoint included in the fit to the target model’s compute. Top Left: Relative error decreases as higher-compute checkpoints are included and then plateaus: for $k\!\in\!\{1,10^{2}\}$ the plateau occurs once the fit includes checkpoints within $\mathord{\sim}$2 orders of magnitude of the target; for $k\!=\!10^{4}$ it occurs within $\mathord{\sim}1.5$ order. Other Three Panels: Backtested estimates of the scaling law parameters converge as the fit includes higher-compute checkpoints, stabilizing to full-fit values once included models are within $\mathord{\sim}2$ orders of magnitude of the target.
We fit the scaling law parameters based on the model checkpoints’ pretraining compute and their corresponding $\mathrm{pass_{\mathcal{B}}}@k$. Visually, the fit scaling laws captured the trend reasonably well for different values of number of attempts per problem $k$ (Fig. 1). We next evaluated what role the number of attempts per problem $k$ has on the scaling law parameters: As $k$ increases, the irreducible error term $E_{0}(k)$ falls roughly exponentially with $k$ and is effectively 0 by $k\approx$1\text{\times}{10}^{2}$\$ (Fig. 2 Left). This is consistent with the probability a problem goes unsolved falling exponentially with the number of attempts per problem (Levi, 2025; Schaeffer et al., 2025b). Thus, by increasing $k$, Eqn. 3 becomes a pure power law with no irreducible error. In comparison, as $k$ increases, the compute prefactor and compute exponent also increase, with the compute prefactor rising $\mathord{\sim}4$ orders of magnitude (Fig. 2 Center) and the compute exponent rising moderately from $1.20\text{\times}{10}^{-1}$ to $3.25\text{\times}{10}^{-1}$ (Fig. 2 Right).
3.2 Predicting Pass Rates from Pretraining Compute
Previously, we fit Eqn. 3 using all available checkpoints. However, a key motivation for scaling laws is prediction: can we forecast the performance of an expensive model using cheaper models?
We evaluated the predictability of pass rates via backtesting, a cross-validation-like process for evaluating forecasts in which cheaper models are used to predict the performance of the most expensive model. We fixed a target model checkpoint pretrained on $C_{\mathrm{target}}$ FLOP; in our work, we used Pythia-12B trained for 300B tokens, pretrained with approximately $2.16\text{\times}{10}^{22}$ FLOP. For a sequence of compute caps $C_{\mathrm{max}}\leq C_{\mathrm{target}}$, we: (i) fit Eq. 3 on all model checkpoints with compute $C\leq C_{\mathrm{max}}$ to obtain $\widehat{E}_{0}(k)$, $\widehat{C}_{0}(k)$, and $\widehat{\alpha}(k)$; (ii) extrapolate to the target compute horizon, and (iii) measure the absolute relative error to the target model’s $\log\!\big(\mathrm{pass}_{\mathcal{B}}@k_{\mathrm{target}}\big)$:
$\mathrm{RelativeError}(k,C_{\mathrm{max}})\operatorname{\stackrel{{\scriptstyle\text{def}}}{{\;=\;}}}\frac{\mathrm{abs}|-\log\!\big(\mathrm{pass}_{\mathcal{B}}@k_{\mathrm{target}}\big)-\widehat{E}_{0}(k)\;-\;\widehat{C}_{0}(k)\cdot C_{\mathrm{target}}^{-\widehat{\alpha}(k)}\big)|}{-\log\!\big(\mathrm{pass}_{\mathcal{B}}@k_{\mathrm{target}}\big)}.$
We report results as a function of the compute ratio $C_{\mathrm{max}}/C_{\mathrm{target}}$.
Relative errors typically decrease and then plateau as higher-compute checkpoints are included in the fit (Fig. 3, top left). For $k\in\{1,1e2\}$, reliable extrapolation requires that the largest checkpoint used for fitting is within $\mathord{\sim}2$ orders of magnitude of the target’s compute, and the relative errors plateau at $1\text{\times}{10}^{-1}$. In comparison, for $k=1e4$, reliable extrapolation requires $\mathord{\sim}1.5$ order of magnitude, and the relative errors plateau below $1$. Similarly, the backtested estimates of the parameters $\widehat{E}_{0}(k),\widehat{C}_{0}(k)$ and $\widehat{\alpha}(k)$ converge to their full-fit values once models within $\mathord{\sim}1.5{-}2.5$ orders of magnitude of $C_{\mathrm{target}}$ are included (Fig. 3, other three panels). For these models on this benchmark, the compute scaling law provides useful forecasts so long as forecasts are made using models within $\mathord{\sim}2$ orders of magnitude of compute as the target model; when only smaller checkpoints are available, the predictive error increases and the fitted parameters are unreliable.
Figure 4: Scaling of Pass Rates with Model Parameters and Pretraining Data (Full Fit). Each panel fits Eqn. 4, $\log\!\big(\mathrm{pass}_{\mathcal{B}}@k\big)(N,D,k)=\mathcal{E}_{0}(k)+N_{0}(k)\,N^{-\beta(k)}+D_{0}(k)\,D^{-\gamma(k)}$. Relative to the compute scaling law, using parameters $N$ and tokens $D$ instead yields tighter in-range fits for all $k$. Consistent with Fig. 2, the irreducible error decreases sharply with $k$ (e.g., ^0: $\widehat{\mathcal{E}}_{0}\!:\ 3.87\!\to\!$0 by $k\!\approx\!10^{2}$), after which variation is dominated by the $(N,D)$ terms. Despite the better global fit, the largest-compute model checkpoint in each panel exhibits comparatively large relative error.
4 Fitting and Predicting Pass Rates from Parameters and Tokens
In our second approach, we assessed how well the benchmark $\mathrm{pass_{\mathcal{B}}@k}$ can be (1) fit and (2) predicted as a function of the number of model parameters $N$ and pretraining tokens $D$.
4.1 Fitting Pass Rates from Model Parameters and Pretraining Tokens
Motivated by Hoffmann et al. (2022b), we posited that the negative log of benchmark pass rates might instead follow a scaling law as a function of model parameters $N$ and pretraining tokens $D$:
$\log\big(\mathrm{pass_{\mathcal{B}}}@k\big)(N,D,k)=\mathcal{E}_{0}(k)+\frac{N_{0}(k)}{N^{\beta(k)}}+\frac{D_{0}(k)}{D^{\gamma(k)}},$ (4)
where $\mathcal{E}_{0}(k)$ is the irreducible error, $N_{0}(k)$ is the parameter scaling prefactor, $\beta(k)$ is the parameter scaling coefficient, $D_{0}(k)$ is the token scaling prefactor and $\gamma(k)$ is the token scaling coefficient.
Figure 5: Predicting Pass Rates from Scaling Parameters and Tokens (Backtesting). We evaluated how accurately the parameters + tokens scaling law (Eqn. 4) predicts the most expensive model’s $\log\!(\mathrm{pass}_{\mathcal{B}}@k)$. Top Left: The relative error decreases as higher-compute checkpoints are included, plateauing once the fit includes models within $\mathrm{\sim}2$ orders of magnitude of the target. Other Five Panels: The estimates of the five scaling law parameters are initially unstable but converge to their full-fit values once included models are within $\mathord{\sim}2$ orders of magnitude of the target.
We fit this parameters + tokens scaling law. Visually, this law provides a better characterization of performance for the full range of models, with the fitted curves appearing tighter and reducing the residual scatter (Fig.4), compared against the pretraining compute law (Fig.1). Consistent with our findings for the compute-only law, the irreducible error term $\widehat{\mathcal{E}}_{0}(k)$ decreases with the number of attempts per problem $k$, although it does so more gradually, reaching zero by $k\approx$3\text{\times}{10}^{2}$\$ (Fig. 10). However, despite the improved global fit, we observed that the largest-compute model checkpoints in our dataset exhibit a comparatively large relative error, suggesting that while this law better explains the performance of less-trained models, it may struggle with extrapolation to the frontier.
4.2 Predicting Pass Rates from Model Parameters and Pretraining Tokens
We evaluated how predictive the parameters + tokens scaling law is using the same backtesting approach (Section 3.2). We again used Pythia 12B model trained for 300B tokens as the target, and iteratively fit Eqn. 4. The relative error in predicting the target model’s performance decreases and then plateaus as more capable models are included in the fit (Fig. 5, Top Left). For all values of $k$, these extrapolations become reliable once the largest checkpoint used for fitting is within approximately $\mathord{\sim}2{-}2.5$ orders of magnitude of the target’s compute. Similarly, the five backtested scaling parameters converge to their full-fit values once models within this same compute range of $\mathord{\sim}1.5-2.5$ orders of magnitude are included (Fig. 5, Other Five Panels). While decomposing pretraining compute into parameters and tokens provides a tighter in-range fit, doing so does not yield a significant improvement in predicting the performance of the most expensive model. For small $k$, the parameters+tokens scaling law has slightly lower relative error at predicting the most expensive model’s performance, but the advantage disappears for larger $k$ (Fig. 5 Upper Left).
Figure 6: Scaling of Pass Rates with Gold Reference Likelihood (Full Fit). Each panel fits Eqn. 6, $\log(\mathrm{pass}_{\mathcal{B}}@k)=\xi_{0}+K_{0}(k)\cdot[-\log\!\big(\mathrm{GoldProb}_{\mathcal{B}}\big)]^{\kappa(k)}$, for $k\in\{1,10^{2},10^{4}\}$. Regressing the gold reference log-likelihoods against the pass rates produces a visually tighter fit with less residual scatter than the compute scaling law (cf. Fig. 1). As $k$ increases, the irreducible error $\widehat{\xi}_{0}(k)$ falls toward zero and the exponent $\widehat{\kappa}(k)$ increases, making the relationship a steeper, purer power law.
5 Fitting and Predicting Pass Rates from Gold Reference Likelihoods
Generative evaluations oftentimes contain not just problems and correct answers, but also “gold references”: high quality responses that reach the correct answer from the given problem. For autoregressive language models, one can straightforwardly compute the likelihood of the gold reference given the problem, as well as the arithmetic mean over the benchmark’s gold reference probabilities:
$\mathrm{GoldProb_{\mathcal{B}}}\quad\operatorname{\stackrel{{\scriptstyle\text{def}}}{{\;=\;}}}\quad\frac{1}{|\mathcal{B}|}\,\sum_{i\in\mathcal{B}}\,p_{\theta}(\text{Gold Reference}_{i}|\text{Problem}_{i}),$ (5)
For our gold reference likelihood scaling law, we assessed how well the benchmark $\mathrm{pass_{\mathcal{B}}@k}$ can be (1) fit and (2) predicted as a function of the log likelihoods of the gold references.
5.1 Fitting Pass Rates from Gold Reference Likelihoods
Motivated by Schaeffer et al. (2024), we tested whether this average likelihood of the gold responses can be accurately regressed to fit and predict benchmark pass rates $\mathrm{pass_{\mathcal{B}}}@k$.
$\log\big(\mathrm{pass_{\mathcal{B}}}@k\big)(\mathrm{GoldProb_{\mathcal{B}}},k)=\xi_{0}(k)+K_{0}(k)\cdot\Big[-\log\big(\mathrm{GoldProb_{\mathcal{B}}}\big)\Big]^{\kappa(k)}$ (6)
Perhaps surprisingly, we discovered that this straightforwardly computable quantity serves as an accurate predictor of benchmark pass rates. As shown in Figure 6, regressing against the negative log likelihood of the gold references produces a visually tighter fit than the pretraining compute law (Fig. 1). The residual scatter is reduced, and the relationship between $\log(\mathrm{pass_{\mathcal{B}}}@k)$ and $\log(\mathrm{GoldProb_{\mathcal{B}}})$ has a smaller reducible error and more closely resembles a power law. We next examined how the number of attempts per problem $k$ influences the scaling law’s parameters (Fig. 11). Consistent with our findings for the prior two scaling laws, the irreducible error term $\xi_{0}(k)$ collapses toward zero as $k$ increases and the exponent $\kappa(k)$ increases smoothly.
Figure 7: Predicting Pass Rates from Gold Reference Likelihoods (Backtesting). While the overall predictive error is comparable to the other laws, it is slightly lower for small $k$ and slightly higher for large $k$. Top Left: The most striking result is the stability of the scaling law’s parameters. In stark contrast to the compute and parameters+tokens laws, whose parameters stabilized only when fit on models within $\mathord{\sim}2$ orders of magnitude of the target, the backtested estimates for $\widehat{\xi}_{0}(k)$, $\widehat{K}_{0}(k)$, and $\widehat{\kappa}(k)$ converge to their final values using models up to $\mathord{\sim}5$ orders of magnitude cheaper than the target. Other Three Panels: This remarkable stability suggests that the relationship between gold reference likelihoods and pass rates provides a robust signal, enabling reliable, long-range forecasts from much cheaper models.
5.2 Predicting Pass Rates from Gold Reference Likelihoods
Using our backtesting approach once more, we evaluated how well the gold references scaling law can predict the performance of our most expensive model. For small $k$, the gold reference likelihoods scaling law has slightly lower relative error at predicting the most expensive model’s performance, but worsens for larger $k$ (Fig. 7 Upper Left). The most remarkable result, however, is the stability of the scaling law’s parameters. In stark contrast to the compute and parameter + token scaling laws whose parameters only stabilized when fit on models within $\mathord{\sim}2$ orders of magnitude of the target, the parameters for the gold reference law are exceptionally robust: the backtested estimates for the three parameters converge to their final values when fitting on models up to $\mathord{\sim}5$ of magnitude cheaper than the target (Fig. 7, Other Three Panels) and for all values of $k$. This significantly stable finding suggests that the relationship between gold reference likelihoods and pass rates is a consistent signal, enabling long-range forecasts from comparatively inexpensive models.
6 How the Compute Scaling Law Emerges from the Parameters + Tokens Scaling Law
The approximately equivalent performance between the compute scaling law (Eqn. 3) and the parameters + tokens scaling law (Eqn. 4) led us to ask whether a deeper connection might exist. We discovered that the compute scaling law is the compute-optimal envelope of the parameters+tokens scaling law. We summarize the insight here and defer a full explanation to Appendix D.
Fix a benchmark $\mathcal{B}$ and attempts per problem $k$. Consider a compute budget $C\approx c\,ND$. The best achievable error is obtained by minimizing the right-hand side of Eqn. 4 over all $(N,D)$ with $ND=C/c$. Evaluating at the optimizer yields Eqn. 3. The mapping of exponents and constants is:
$\alpha(k)=\Bigl(\tfrac{1}{\beta(k)}+\tfrac{1}{\gamma(k)}\Bigr)^{-1},\qquad E_{0}(k)=\mathcal{E}_{0}(k),$
with a closed-form $C_{0}(k)$.
Compute-Optimal Allocation. Along the envelope (i.e., when Eqn. 3 is tight), the optimal split of compute between parameters and tokens is
$N^{*}(C,k)=\Bigl(\tfrac{\beta\,N_{0}(k)}{\gamma\,D_{0}(k)}\Bigr)^{\!\frac{1}{\beta+\gamma}}\Bigl(\tfrac{C}{c}\Bigr)^{\frac{\gamma}{\beta+\gamma}},\qquad D^{*}(C,k)=\Bigl(\tfrac{\gamma\,D_{0}(k)}{\beta\,N_{0}(k)}\Bigr)^{\!\frac{1}{\beta+\gamma}}\Bigl(\tfrac{C}{c}\Bigr)^{\frac{\beta}{\beta+\gamma}}.$
Thus the parameter/token ratio shift with scale according to $\beta,\gamma$; when $\beta=\gamma$, both grow like $C^{1/2}$.
Departures from the Envelope. If training does not follow the compute-optimal allocation $(N^{*},D^{*})$, define the misallocation ratio $r(C)=\tfrac{N}{N^{*}(C)}=\tfrac{D^{*}(C)}{D}$. The departure from the envelope is multiplicative on top of Eqn. 3:
$\log\big(\operatorname{pass}_{\mathcal{B}}\text{@}k\big)(C,k)=E_{0}(k)+C_{0}(k)\cdot C^{-\alpha(k)}\cdot\underbrace{\Bigl[\tfrac{\gamma}{\beta+\gamma}\,r^{-\beta}+\tfrac{\beta}{\beta+\gamma}\,r^{\gamma}\Bigr]}_{\displaystyle\Phi(r;\beta,\gamma)\ \geq 1},$
with equality iff $r\equiv 1$. Small deviations are benign ($\Phi=1+\tfrac{\beta\gamma}{2}\,(\log r)^{2}+O(\log r)^{3})$ ), but persistent off-ridge scaling yields a growing penalty. A useful case: if one scales $N$ and $D$ at a fixed ratio (i.e., without re-optimizing with $C$), the effective compute slope degrades to $\alpha_{\text{path}}=\frac{\min\{\beta,\gamma\}}{2}$.
Takeaways.
(i) The compute law is not an independent phenomenology; it is the optimal-allocation shadow of the parameters + tokens law. (ii) Use the allocation above to stay on-envelope when planning larger models. (iii) Implementation constants relating $C$ to $ND$ shift $C_{0}(k)$, but do not change $\alpha(k)$. All statements hold pointwise in $k$.
7 Discussion
In this work, we investigated pretraining scaling laws for generative evaluations of language models. Our work was motivated most directly by OpenAI et al. (2024)’s success at predicting pass rates at HumanEval (Chen et al., 2021), as well as by Hu et al. (2024).
Key Findings Our work’s main contribution is the proposal and rigorous backtesting of three distinct scaling laws for pass-at-$k$, using (1) pretraining compute, (2) model parameters and pretraining tokens, and (3) gold reference log likelihoods as covariates. We demonstrate that a key hyperparameter in generative evaluation – the number of attempts per problem $k$ – offers a novel lever for controlling the scaling law parameters and predictability. Empirically, we find that while all three laws have comparable predictive accuracy, the parameters of the gold reference likelihood law are uniquely stable, converging across nearly five orders of magnitude of compute. Finally, we establish a theoretical connection between the compute law and the parameters + tokens law, proving that the compute law emerges as the compute-optimal envelope of the parameters-and-tokens law.
Limitations The main limitation of this work is that the empirical results are based on a single benchmark (GSM8K) and a single model family (Pythia). Additionally, we consider only pass-at-$k$ generative evaluations, whereas other are of course worth considering as well.
Related Work Due to space constraints, we defer related work to Appendix A.
Future Directions Multiple next steps are clear. First, subsequent work should determine whether predictions can be made more accurately using either additional sampling or controlled pretrained models. Second, OpenAI et al. (2024) claimed subsetting and/or stratifying problems enabled better predictions; what exactly did this approach entail, and how much did it help? Third, the utility of the gold reference log likelihoods was perhaps surprising because if we think of next-token sampling as a branching process, there are likely exponentially many paths from the problem to the correct answer, and it is unclear that the benchmark-provided gold references have any meaningful relationship with likely samples from the model; this raises at least two future questions: (i) how general is this relationship across benchmarks and models, and (ii) to what extent can this indicator be trusted for predicting the performance of more expensive models, especially under optimization pressure?
Appendix A Related Work
The study of neural scaling laws, which characterize how model performance predictably improves with resources, has become a cornerstone of modern large-scale machine learning (Hestness et al., 2017; Kaplan et al., 2020a; Hoffmann et al., 2022a; Roberts et al., 2022; Bahri et al., 2024). Foundational work demonstrated that pretraining loss follows a power law with respect to model size, dataset size, and training compute (Kaplan et al., 2020a; Henighan et al., 2020; Hoffmann et al., 2022a). These principles now guide the development of nearly all state-of-the-art language models (OpenAI et al., 2024; Team et al., 2024; Grattafiori et al., 2024; DeepSeek-AI et al., 2025; Qwen et al., 2025; Zhipu-AI et al., 2025). This research area has since expanded to explore scaling properties in diverse domains, including computer vision (Zhai et al., 2022), machine translation (Ghorbani et al., 2021), and even for optimizing inference-time compute rather than training (Wu et al., 2024; Snell et al., 2024). Despite this breadth, scaling laws for complex, multi-step generative tasks—which often represent the ultimate goal of language modeling—have remained comparatively underexplored due to the challenges of evaluation. Our work builds most directly on recent efforts to forecast performance on such generative evaluations. The GPT-4 Technical Report pioneered this direction by showing that the negative log pass rate on a subset of the HumanEval coding benchmark scaled predictably as a power law of pretraining compute (OpenAI et al., 2024). They successfully predicted GPT-4’s performance by extrapolating from smaller models over approximately three orders of magnitude (OpenAI et al., 2024). We adopt their proposed functional form for our compute-based law and extend their work by systematically analyzing how the number of attempts, $k$, reshapes the law’s parameters. Furthermore, our rigorous backtesting reveals the brittleness of this compute-only law, finding it reliable only across 1-2 orders of magnitude in our setting. Separately, our work shares a philosophical motivation with Hu et al. (2024), who investigate emergent abilities. They argue that seemingly unpredictable jumps in performance on discriminative tasks can be understood as smooth, continuous improvements when measured by the model’s underlying log-likelihoods. We apply a similar insight to a different problem: forecasting. We operationalize the log probability of gold-standard solutions not just as an explanatory tool, but as a direct covariate in a predictive scaling law for generative tasks, and we empirically demonstrate its superior stability for long-range extrapolation.
Appendix B Methodology: Number of Samples Per Problem Per Model
Because sampling from language models is expensive and because our compute budget is limited, we needed to judiciously choose how many samples to draw per problem per model. A key consideration is that the number of samples per problem sets the resolution, i.e., if we draw $n$ samples per problem, then any pass rate on that problem below $1/n$ will likely appear to be 0. Since pass rates typically increase with pretraining compute, we chose a loose heuristic based on the pretraining compute of each model. For a model with $C$ pretraining FLOP, we aimed to sample
$\text{Num. Samples per Problem }=\operatorname{clip}(100\cdot 10^{23-\log_{10}(C)},3.2e4,1e6).$
$3.2e4$ was chosen so that we would have more samples per problem than the largest number of attempts per problem $k$ we considered $(1e4)$. Fig. 8 shows how many samples per problem per model we had at the time our analyses were conducted. Additional sampling would have likely benefited our analyses.
Figure 8: Number of Samples per Model per Problem. This heatmap visualizes how many samples we had drawn per model (y-axis) per problem (x-axis) at the time our analyses were conducted.
Appendix C Scaling Law Parameters by Number of Attempts Per Problem $k$
For the three scaling laws we consider, we visualize how the fit parameters change as a function of the number of attempts per problem $k$ (Fig. 9, Fig. 10, Fig. 11).
Figure 9: Pretraining Compute Scaling Law Parameters by Number of Attempts per Problem (Full Fits). We visualize how the parameters of Eqn. 3 change with respect to $k$. Hue corresponds to $k$.
Figure 10: Model Parameters and Tokens Scaling Law Parameters by Number of Attempts per Problem (Full Fits). We visualize how the parameters of Eqn. 4 change with respect to $k$. Hue corresponds to $k$.
Figure 11: Gold Reference Log Likelihoods Scaling Law Parameters by Number of Attempts per Problem (Full Fits). We visualize how the parameters of Eqn. 5 change with respect to $k$. Hue corresponds to $k$.
Appendix D Relation Between the Compute Scaling Law and the Parameters + Tokens Scaling Law
Here, we provide a concrete analytic relation between Eqn. 3 and Eqn. 4 provided in the main text. For a fixed benchmark $\mathcal{B}$ and number of attempts per problem $k$, we model
$-\log\!\bigl(\operatorname{pass}_{\mathcal{B}}\text{@}k\bigr)(N,D,k)=\mathcal{E}_{0}(k)+\frac{N_{0}(k)}{N^{\beta(k)}}+\frac{D_{0}(k)}{D^{\gamma(k)}},$ (7)
and, when training under a compute budget C=cNDC=c\,ND,
$-\log\!\bigl(\operatorname{pass}_{\mathcal{B}}\text{@}k\bigr)(C,k)=E(k)+\frac{C_{0}(k)}{C^{\alpha(k)}}.$ (8)
Compute-optimal envelope.
Assume the excess loss beyond the irreducible term is additive and separable in $(N,D)$ as in Eq. 7, with $N_{0}(k),D_{0}(k)>0$ and exponents $\beta(k),\gamma(k)>0$. For each fixed $k$, define
$A=N_{0}(k),\quad B=D_{0}(k),\quad\beta=\beta(k),\quad\gamma=\gamma(k),$ (9)
and
$F(N,D)=\frac{A}{N^{\beta}}+\frac{B}{D^{\gamma}},\qquad\mathcal{F}(C)=\min_{\begin{subarray}{c}N,D>0:\;\\ cND=C\end{subarray}}F(N,D).$ (10)
Then the compute law Eq. 8 is the compute-optimal envelope of Eq. 7, with
$E(k)=\mathcal{E}_{0}(k),\qquad\alpha(k)=\frac{\beta(k)\,\gamma(k)}{\beta(k)+\gamma(k)}=\Bigl(\tfrac{1}{\beta(k)}+\tfrac{1}{\gamma(k)}\Bigr)^{-1},$ (11)
$C_{0}(k)=(\beta+\gamma)\,\beta^{-\frac{\beta}{\beta+\gamma}}\,\gamma^{-\frac{\gamma}{\beta+\gamma}}\,A^{\frac{\gamma}{\beta+\gamma}}\,B^{\frac{\beta}{\beta+\gamma}}\,c^{\alpha(k)}.$ (12)
The compute-optimal allocation along the envelope is
$N^{*}(C,k)=\Bigl(\frac{\beta A}{\gamma B}\Bigr)^{\!\frac{1}{\beta+\gamma}}\Bigl(\frac{C}{c}\Bigr)^{\frac{\gamma}{\beta+\gamma}},\qquad D^{*}(C,k)=\Bigl(\frac{\gamma B}{\beta A}\Bigr)^{\!\frac{1}{\beta+\gamma}}\Bigl(\frac{C}{c}\Bigr)^{\frac{\beta}{\beta+\gamma}}.$ (13)
Derivation Sketch.
Form the Lagrangian $\mathcal{L}=AN^{-\beta}+BD^{-\gamma}+\lambda\,(ND-C/c)$. First-order conditions give $\beta AN^{-\beta}=\gamma BD^{-\gamma}$ and $ND=C/c$, from which (A.5) follows. Substituting $N^{*},D^{*}$ into $F(N,D)$ yields $\mathcal{F}(C)=C_{0}(k)\,C^{-\alpha(k)}$ with $\alpha(k)$ and $C_{0}(k)$ as in Eqs. 11-12, and $E(k)=\mathcal{E}_{0}(k)$.
Remarks.
  1. 1.
    The compute exponent is the parallel sum of the $(N,D)$ exponents: $\alpha=(1/\beta+1/\gamma)^{-1}$.
  2. 2.
    If $\beta=\gamma$, then $\alpha=\beta/2$ and $N^{*},D^{*}\propto C^{1/2}$ with $C_{0}=2\,\beta^{-1}(AB)^{1/2}c^{\beta/2}$.
  3. 3.
    If one scales $N$ and $D$ at a fixed ratio (i.e., without re-optimizing with $C$), the effective compute exponent degrades to $\alpha_{\text{path}}=\min\{\beta,\gamma\}/2$.
  4. 4.
    All quantities may depend on $k$; the mapping in Eqs. 11-13 applies point-wise in $k$. The conversion constant $c$ enters only via $C_{0}$, not $\alpha$.
D.1 Multiplicative deviation from the optimal compute envelope
Let $(N^{*}(C),D^{*}(C))$ denote the compute–optimal allocation in Eq. 13 and define the misallocation ratio
$r(C)\coloneqq\frac{N(C)}{N^{*}(C)}=\frac{D^{*}(C)}{D(C)}.$ (14)
A direct algebraic manipulation using the first-order optimality condition $\beta A(N^{*})^{-\beta}=\gamma B(D^{*})^{-\gamma}$ gives the exact factorization
$\frac{-\log\!\bigl(\operatorname{pass}_{\mathcal{B}}\text{@}k\bigr)(C)-\mathcal{E}_{0}}{-\log\!\bigl(\operatorname{pass}_{\mathcal{B}}\text{@}k\bigr)_{\text{opt}}(C)-\mathcal{E}_{0}}=\Phi(r;\beta,\gamma)\coloneqq\frac{\gamma}{\beta+\gamma}\,r^{-\beta}+\frac{\beta}{\beta+\gamma}\,r^{\gamma}\;\;\geq 1,$ (15)
with equality iff $r\equiv 1$. Equivalently, this produces a multiplicative correction to Eq. 8
$-\log(\operatorname{pass}_{\mathcal{B}}\text{@}k)(C)=\mathcal{E}_{0}+C_{0}\,C^{-\alpha}\,\underbrace{\Phi\!\left(\tfrac{N}{N^{*}};\beta,\gamma\right)}_{\text{misallocation factor}\;\geq 1}.$ (16)
Thus the deviation from equality is fully parameterized by $r$ via the dimensionless penalty $\Phi\geq 1$.
Local and asymptotic behavior of the penalty.
Let $t=\log r$. A Taylor expansion of equation 15 around $r=1$ gives
$\Phi(r;\beta,\gamma)=1+\frac{\beta\gamma}{2}\,t^{2}+O(t^{3}),$ (17)
so small misallocations incur only a quadratic penalty in $\log$-space. Along a power-law path $N\propto C^{p}$ (with constants absorbed into $r$), $t=\bigl(p-\tfrac{\gamma}{\beta+\gamma}\bigr)\log C+O(1)$; hence if $p\neq\tfrac{\gamma}{\beta+\gamma}$ the penalty grows polynomially:
$\Phi(r(C))\asymp\begin{cases}\frac{\beta}{\beta+\gamma}\,r(C)^{\gamma},&r(C)\to\infty\quad(\text{over-allocating }N),\\ \frac{\gamma}{\beta+\gamma}\,r(C)^{-\beta},&r(C)\to 0\quad(\text{under-allocating }N).\end{cases}$ (18)
This isolates the compute–optimal scaling $C^{-\alpha}$ and expresses all off-ridge effects through a single, dimensionless factor depending only on the allocation ratio $N/N^{*}$ (or equivalently $D^{*}/D$).
We conclude that if one matches the compute law without optimizing $(N,D)$, the result is not a single power law in general: it is a compute–optimal power multiplied by a misallocation penalty $\Phi\!\left(N/N^{*}\right)\geq 1$.